OpenStudy (imnotgoodatmath):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 29 vehicles and 82 tires. How many motorcycles passed his house? 6 10 12 17

OpenStudy (anonymous):

Let m= motorcycles Let = cars m+c= 29 vehicles -------------------------- a car has 4 wheels, and a motorcyle has 2 wheels. so 4c + 2m = 82 tires --------------------------

OpenStudy (anonymous):

so we have m+c = 29 vehicles c= 29-m Replace this value of c in the 2nd equation. 4(29-m) +2m= 82 116 -4m+2m= 82 34 = 2m m= 17 so there were 17 motorcycles.